# Consumption Smoothing in Metropolis: Evidence from the Working-class Households in Prewar Tokyo

### 26 Jan 2024 v3

@@ A.5 Machinery Factories in Tsukishima
First, I use the results of the Elementary School Children Survey included in the Tsukishima Survey to analyze the occupation (industry) of the head of the family household in Tsukishima. This is a survey of 4th-6th graders at Tsukishima 1st and 2nd Elementary Schools in July 1919. There were 1,173 children in the 3rd grade, and the survey included all 1,100 children who resided in Tsukishima. At that time, there were no other elementary schools besides those mentioned above. Thus, this survey is a complete survey for all children in grades 4-6 in Tsukishima. Although it does not cover households without children in grades 4-6, it is still a beneficial survey that provides an overview of the family households this paper covers. Panel A of Table A.5 shows the industrial structures. The composition ratios are similar to the census-based statistics shown in Panel A of Table 1. It does not specify whether brothers and sisters in the same household were counted twice. Despite this, this result confirms that the statistics obtained from this survey are useful in identifying overall trends in family households in Tsukishima.
Panel B of Table A.5 divides the “social classes” of the household heads in the industrial sector into six categories. The report considers large business owners to be the capitalist class, small business owners, free enterprise owners, directors to be the middle class, and workers to be the working class (Department of Health, Ministry of the Interior 1923a, p. 50). The table reveals at least three things. First, most heads of households in the industrial sector are workers, with only 1% of the capitalist class and less than 20% of the middle class. Although the definition of “executive” is unclear, it is unlikely that “small business owners” are households with a reasonably large income difference from skilled worker households because they must be small factory owners (Department of Health, Ministry of the Interior 1923a, p. 64).
Table A.5: Industrial Structure and Social Classes in Tsukishima
@@
Notes: This table summarizes the statistics from the Complete Survey on the 4th-6th graders in Tsukishima1st and 2nd Elementary Schools in July 1919. Panel A summarizes the heads’ industries. Panel B sorted the “social classes” of the heads in the manufacturing sector. The report does not specify the definition of each “social class”. Source: Department of Health, Ministry of the Interior 1923a, pp. 386–387.
@@
Panel B of Table A.6 classifies the types of the 620 workers based on the results of the primary school survey used in Table A.5. There were 455 skilled and 165 unskilled workers. Notably, these 620 workers include those not in the industrial sector: 490 industrial workers, 13 commercial, 61 in transportation, 16 in public service free trade, 38 in other gainful employment, and 2 in agriculture and mining in total. Since the workers in the industrial sector should include factory workers, workers in the cottage industry, and handicraft workers, the day laborers are considered to be classified into non-industrial workers (in fact, $620$ workers $-120$ day laborers $=490$). Therefore, it is estimated that roughly 90% of factory workers ($455/490$) were skilled workers, $65$% of whom worked in factories, $17$% in cottage industries, and $18$% as craftsmen. Since the workers in cottage and handicraft industries are considered to work in small factories, $65$% of these factory workers might had worked in the middle- to large-scale factories (with 15+ worerks).
One must be careful about the fact that this survey only covers the heads who had primary school children aged 9–11 (i.e., 4th to 6th grades), suggesting that most of these heads are in their 30s. Therefore, the share of skilled workers to the entire workers should be smaller than the $90$%. At least, however, it is plausible to see that almost all of the Tsukishima workers in their 30s were skilled workers at that time.
@@
Panel A: Scale of Factories
@@
Survey subject All factories
Machinery and Drum Casting Smithing Total
@@
Notes: Panel A summarizes the machinery factories by the scale of the factory. Statistics are from the manufacturing census conducted in the Tsukishima Survey (1st November 1920). The threshold (i.e., 15 workers) is based on the application of the Factory Act enacted in 1916. Panel B sorted the heads by their skill level and job type. Statistics are obtained from the Complete Survey on the 4th-6th graders in Tsukishima1st and 2nd Elementary Schools in July 1919. Source: Panel A: Department of Health, Ministry of the Interior 1923a, pp. 386–387; 410–414. Panel B: Department of Health, Ministry of the Interior 1923a, p. 149.
@@ A.6 Works in the Machinery Factories
39% of the factories in Tsukishima are machinery and equipment manufacturing factories, the largest share of the machinery industry. Typical jobs for workers in machinery and equipment manufacturing plants are finishers (or fitters) and turners. A finisher is “a skilled worker responsible for finishing, assembling, and commissioning machine products” and is “one of the finest skilled workers because it requires skills in all aspects to correct defects in machine making” (p. 393). The turner is the “middle axis of machinists,” to which most semi-skilled and skilled workers belong. Becoming a full-fledged finisher and lathe operator takes at least five years. Generally, apprentices enter the workforce around 14–18, but they do not work in the same factory for many years within the apprenticeship system. After learning the skills for 1-2 years, they often move to another factory to improve their skills (p. 394).
Wages are paid on a piece-work basis or on a time basis. In Tsukishima, a time contract was adopted, and payment was based on the “residual time” method. Under the “residual time” method, “if the work is completed within the contracted time, the remaining time is considered the bonus.” For example, if 12 hours are contracted, and the work is completed in 10 hours, 2 hours are paid as a bonus (pp. 396-397). Wages were paid by dividing one month into two periods (first and second semester). The average workday in a machine shop is reported to be roughly 10 hours, with occasional all-night work (p. 402). Overtime hours were reportedly around 2 hours (p. 402).
@@
Next, I summarize labor in the canning factory, which accounts for 24% of all Tsukishima factories. The can factory is said to be the “representative industry of Tsukishima”. It is said to be engaged in shipbuilding, bridges, steel frames for construction, and other works, including fabrication and repair of gas reservoirs (p. 406). Workers in a canning factory are called “boiler makers” and are typically involved in bending steel plates and joining them with rivets. Workers are divided into rivetter, holder-up, plater, angle-ironsmith, and caulking. The work is done in groups, with the skilled angle-ironsmith and caulking. The “bōsin” of the rivetter gives the instructions as the head of the group. Wages are paid on a piece-rate contract basis. Because of the contracted wage, the cannery workers’ income is unstable; thus, they are paid a fixed wage (daily wage amount) during the off-season.
@@
Smithing is the second most common industry, accounting for 28% of the total. Panel A of Table A.6 reveals that most of the smithing industry is relatively small scale. They make boulders, nuts, rivets, lathe stops, jigs, and spanners for can manufacturing and construction. They were subcontractors who manufactured metal parts to supply the large factories on the island (p. 412). There were three types of jobs: yokoza, sakite, and tatara-fuki (bellows-blower). The yokoza were self-employed skilled workers (oyakata), while the other two were young unskilled laborers. In small operations, wages were based on a percentage based on voluntary business discussions with the proprietor, and working hours were not regular (pp. 411–413).
@@ B.1 Tsukishima Survey: The First Social Survey in Japan
In modern Japan, the methodology of social research was formed in the 1930s. Before this, the generation of social research methodologies was attempted in the 1910s and 1920s (Kawai 1980, pp. 47–48). Kawai (1980, p.48) summarizes the history of social research methodology: during this period, “numerous social surveys were conducted by various survey entities, including government agencies, research groups, local administrative bodies, private organizations, survey research institutes, and researchers.” This is because “around WWI, there were many problems in labor, lifestyle, urban, health and sanitation, public safety, and civic due to the heavy and chemical industrialization, the monopolization of capitalism, and urban expansion”. The “Report on Field Survey in Tsukishima, Kyobashi Ward, Tokyo” (called Tsukishima Survey) used in this paper is the most pioneering of these surveys of urban communities. [^73] Takano himself, who was responsible for overseeing the survey, refers to this field survey as the “Tsukishima Survey” (Department of Health, Ministry of the Interior 1923a, p. 6). Considering this, I will refer to it as the Tsukishima Survey throughout this paper. The Tsukishima Survey was a pioneering household survey in Japan (Sekiya 1970, p. 43).
@@
First, I summarize the investigators, the research questions, and the subjects of the Tsukishima Survey. The person responsible for the Tsukishima Survey was Iwasaburō Takano (hereafter “Takano”), a professor of social statistics and social policy at the Law School of Tokyo Imperial University. Takano became a professor of economics at Tokyo Imperial University in 1919, and he became the director of the OISR in 1920. Takano placed great importance on gathering objective statistics about workers through social surveys (Sekiya 1970, p. 33). He had the experience of conducting the first household survey in Japan in 1916.747474This is called the “Survey of Household Income and Expenditure of 20 Artisans and Workers in Tokyo” and was inspired by Charles Booth (Miyoshi 1980, p. 33). Although the original documents of this survey were preserved at Tokyo Imperial University, those were destroyed in the Great Kantō Earthquake (Takano 1933, pp. 727–728). More precisely, before the 1916 household survey, Takano conducted a field survey in a poor area of Tokyo while a student at Tokyo Imperial University reported his findings in an April 1894 exercise entitled “East London in Tokyo” (Kawai 1981b, p. 3). Takano has been a member of the seventh subcommittee (Rural Sanitation Subcommittee) of the Ministry of Home Affairs’ Health and Sanitation Investigation Committee since 1916.757575The Health and Sanitation Investigation Committee of the Ministry of Home Affairs was established on June 27, 1916, under the second Shigenobu Okuma Cabinet by issuing Imperial Ordinance No. 172. Initially, 34 members were appointed, and the first meeting was held on July 8. The reason for the establishment of this committee was the need for a policy response regarding the establishment and institutionalization of various investigation organizations in the establishment of the national system and the penetration of the capitalist economy during the Meiji period. Another reason is the requirement for a public health response to the rising mortality rate, exemplified by the tuberculosis epidemic (Kawai 1981a, pp. 10–13). In 1917, the Seventh Section evolved into the Section on Urban and Rural Sanitary Conditions (Kawai 1981a, pp. 12–13). On October 22, 1918, Takano submitted a proposal to the subcommittee for a field survey of urban sanitary conditions, which was approved. The survey aimed to determine the housing conditions, household budget status, and health status (child health, production, stillbirths, deaths, and diseases) in the areas where skilled-worker families congregated (Sekiya 1970, p. 6). Initially, the survey area was planned to be Yanagibashi Yokogawa-chō in Honjo Ward. However, at a committee meeting on November 23, the survey site was changed to Tsukishima in Kyobashi Ward, which had more skilled worker families than in Yokogawa-cho (Kawai 1980, pp. 55–59). [^76] Takeda (2015, pp. 75–76) provides more details on how the survey site was changed. Takano was looking for a commissioned researcher to conduct the survey and asked Yoshino Sakuzō, a colleague at Tokyo Imperial University, to recommend a suitable person. Yoshino approached his former student Kotora Tanahashi, who had become a member of the fraternity club, a mutual aid organization for workers, to accept the investigator position. The diary written by Tanahashi Kotora (“Tanahashi Kotora Diary”) is in the possession of the OISR, and Takeda used the diary as the basis for her analysis of the circumstances surrounding the change of survey location. Tanahashi eventually recommended Yoshitsuru Yamana, who was his classmate in high school and Tokyo Imperial University, as a researcher and considered Tsukishima to be the best research area because of the ease of selecting survey households using the fraternity’s network (Takeda 2015, pp. 76–77). On October 26, Tanahashi and Yamana, together with the Kyobashi Branch of the fraternity club, conducted an on-site inspection of Tsukishima. Based on this inspection, Tanahashi and Yamana approached Takano about locating the survey station in Tsukishima. Finally, Takano and Yamana conducted an on-site inspection of Tsukishima on October 30 and decided on Tsukishima as the survey site.
@@
While the statistical method of sampling was not yet widespread at that time, Takano selected the survey sites based on his visionary thinking. In fact, “There can be no dispute that we have chosen Tokyo as the target of our metropolitan survey. However, it is impossible to survey the entire Tokyo area. We have no choice but to conduct a partial survey. For this purpose, it would be the best idea to select a representative area in Tokyo and survey that area narrowly but deeply.” (Department of Health, Ministry of the Interior 1923a, pp. 2–3). Takano further explained the need to focus on the working-class, which makes up most of the citizens, rather than the political and commercial center of Tokyo: “Most of Tokyo’s citizens are wage earners who subsist on labor. As is usual in large cities, a large number of people are engaged in manual labor in Tokyo. However, the most genuine type of worker is the skilled worker. Therefore, the area where many families of skilled workers congregate can be regarded as a representative area of Tokyo. And it seemed to me that a study of the social and sanitary conditions in that area could explain a large part of the social and sanitary conditions in Tokyo.” (Department of Health, Ministry of the Interior 1923a, pp. 3).
@@
It is necessary to delve into the point where Takano states that “the most genuine type of worker is the skilled workers” and that the main target of the Tsukishima Survey was skilled workers. [^77] In other words, the Tsukishima Survey was not designed for poor households. Therefore, unlike the surveys conducted in the UK by C. Booth and B.S. Rowntree, no standard of living was set in the survey (Sekiya 1970, p. 42). The description by Kawai (1981a, pp. 13–18) is helpful in this regard. As summarized in Section 2, Japan’s heavy and chemical industrialization throughout WWI led to a large concentration of city workers. As a result, the ratio of small-scale conventional factories based on “master employees” as is often observed in the textile industry decreased. On the other hand, the number of factories with a direct “factory owner-employee” employment relationship increased, particularly in the machinery and equipment, metal, and chemical industries. These workers in the modern industrial sector accumulated skills, formed households, and settled in the cities. It can be said that a new urban worker lifestyle was being formed. In short, although the city had a class of workers engaged in traditional industries, a new type of workers in the modern industrial sector, and a lower class, Takano believed that the skilled workers in the modern industrial sector would play a central role in terms of the labor force, labor movement, and urban life.
@@
In addition, Sekiya notes the rise of the labor movement. As noted in Section 2, while wages rose throughout WWI, the price surge caused relative wages to decline. In 1916, Japan’s first labor protection legislation, the Factory Law, was enacted. However, the number of working days and hours did not decrease, and the number of labor disputes steadily increased (Sekiya 1970, pp. 24–26). Given this, Takano believed that workers should organize labor unions to help themselves with labor problems that could not be resolved through labor protection legislation. Sekiya (1970, pp. 26–40) speculates that Takano’s intention in conducting the Tsukishima Survey was to obtain basic data for solving labor problems, specifically to identify the scope of matters to be solved by workers on their own initiative (or covered by labor protection legislation). His insights suggest that this historical background must be addressed as a reason for Takano’s focus on skilled workers.
@@
Kawai (1981b, pp. 15; 17) assessed the 40 households covered by the household survey included in the Tsukishima Survey as “the representativeness of the sample has remained ambiguous.” There is some mention but no clear description of how the sample households were characterized among the worker households in Tsukishima. It is reported that “to find suitable persons to fill in the form, we must ask for the assistance of school principals, police officers, physicians, and laborers, and we have requested a meeting of these persons at the survey office to explain the purpose of the survey, and to solicit applicants for the distribution of the household account book.” (Department of Health, Ministry of the Interior 1923a, pp. 8–9). As explained, the target households were not randomly selected because the statistical method of sampling was not prevalent (Section 3 provides a detailed analysis of the characteristics of the surveyed households). However, it is plausible that Takano designed this household budget survey in the sampling framework to capture the representative households in Tsukishima.

### 25 Feb 2024 v6

@@ A.5 Machinery Factories in Tsukishima
First, I use the results of the Primary School Children Survey included in the Tsukishima Survey to analyze the occupation (industry) of the household heads in Tsukishima. This is a survey of 4th-6th graders at Tsukishima 1st and 2nd Elementary Schools in July 1919. There were 1,173 children in these grades, and the survey included all 1,100 children who resided in Tsukishima. At that time, there were no other elementary schools besides those mentioned above. Thus, this survey is a complete survey for all children in grades 4-6 in Tsukishima. Although it does not cover households without children in grades 4-6, it is undoubtedly a beneficial survey that provides an overview of the households with children that are the main target of this study.
Panel A of Table A.5 shows the industrial structures. The composition ratios are similar to the census-based statistics reported in Panel A of Table 1. It does not specify whether brothers and sisters in the same household were counted twice. Despite this, this similarity confirms that the statistics obtained from this survey are useful in identifying overall trends in family households in Tsukishima.
Panel B of Table A.5 divides the “social classes” of the household heads working in the manufacturing sector into six categories. In the report, large business owners are classified into the capitalist class; small business owners, free enterprise owners, and directors are regarded as the middle class; workers are defined as the working class (Department of Health, Ministry of the Interior 1923a, p. 50). The table reveals at least three things. First, most heads of households are workers, with only 1% of the capitalist class and less than 20% of the middle class. While the definition of “executive” is unclear, “small business owners” must be small factory owners who could have similar earnings to the skilled worker households (Department of Health, Ministry of the Interior 1923a, p. 64).
Table A.5: Industrial Structure and Social Classes in Tsukishima: Evidence from Tsukishima Primary School Survey
@@
Notes: This table summarizes the statistics from the complete survey on the 4th-6th graders in Tsukishima1st and 2nd Primary Schools in July 1919. Panel A summarizes the heads’ industries, and Panel B sorts the “social classes” of the heads in the manufacturing sector, respectively. The report does not specify the definition of each “social class”. Source: Department of Health, Ministry of the Interior 1923a, pp. 386–387.
@@
Panel B of Table A.6 classifies the types of the 620 workers based on the results of the primary school survey used in Table A.5. There were 455 skilled and 165 unskilled workers. Notably, these 620 workers include non-industrial workers: 490 industrial workers, 13 commercial, 61 in transportation, 16 in public service free trade, 38 in other gainful employment, and 2 in agriculture and mining in total. Since the workers in the industrial sector should include factory workers, workers in the cottage industry, and handicraft workers, the day laborers are considered to be classified into non-industrial workers (in fact, $620$ workers $-120$ day laborers $=490$). Therefore, roughly 90% of factory workers ($455/490$) can be regarded as skilled workers, $65$% of whom worked in factories, $17$% in cottage industries, and $18$% as craftsmen. Since the workers in cottage and handicraft industries are considered to work in small factories, $65$% of these workers might have worked in the middle- to large-scale factories (with 15+ workers).
One must be careful that this survey only covers the heads with primary school children aged 9–11 (i.e., 4th to 6th grades), suggesting that most of these heads are in their 30s. Therefore, the share of skilled workers to the entire workers should be smaller than the $90$%. At least, however, it is plausible to see that most of the Tsukishima workers in their 30s were skilled workers at that time.
@@
Panel A: Scale of Machinery Factories
@@
Survey subject All machinery factories
Machinery and Can Casting Smithing Total
@@
Notes: Panel A summarizes the number of machinery factories by the scale and type of the factory. Statistics are from the manufacturing census conducted in the Tsukishima Survey (1st November 1920). The threshold of the scale (i.e., 15 workers) is based on the application of the Factory Act enacted in 1916. Panel B sorted the heads by their skill level and job type. Statistics are obtained from the Complete Survey on the 4th-6th graders in Tsukishima1st and 2nd Primary Schools in July 1919. Source: Panel A: Department of Health, Ministry of the Interior 1923a, pp. 386–387; 410–414. Panel B: Department of Health, Ministry of the Interior 1923a, p. 149.
@@ A.6 Works in the Machinery Factories
Tsukishima Survey includes the factory survey that counts the number of machinery factories in Tsukishima. Panel A of Table A.6 summarizes the number of machinery factories by type of plant. The machinery and equipment manufacturing factories shared 39% of all the machinery factories. Typical jobs for workers in machinery and equipment manufacturing plants are finishers (or fitters) and turners. A finisher is “a skilled worker responsible for finishing, assembling, and commissioning machine products” and is “one of the finest skilled workers because it requires skills in all aspects to correct defects in machine making” (Department of Health, Ministry of the Interior 1923a, p. 393). The turner is the “middle axis of machinists,” to which most semi-skilled and skilled workers belong. Becoming a full-fledged finisher and lathe operator takes at least five years. Generally, apprentices enter the workforce around 14–18, but they do not work in the same factory for many years within the apprenticeship system. After learning the skills for 1-2 years, they often move to another factory to improve their skills (Department of Health, Ministry of the Interior 1923a, p. 394). Wages are paid on a piece-work basis or on a time basis. In Tsukishima, a time contract was adopted, and payment was based on the “residual time” method. Under the “residual time” method, “if the work is completed within the contracted time, the remaining time is considered the bonus.” For example, if 12 hours are contracted, and the work is completed in 10 hours, 2 hours are paid as a bonus (pp. 396-397). Wages were paid by dividing one month into two periods (first and second semester). The average work hours are reported to be roughly 10 hours, with occasional all-night work (p. 402). Overtime hours were reportedly around 2 hours per day (Department of Health, Ministry of the Interior 1923a, p. 402).
@@
Next, I summarize labor in the canning factory, which accounts for 24% of all Tsukishima factories. The canning factory is engaged in shipbuilding, bridges, steel frames for construction, and other works, including fabrication and repair of gas reservoirs (p. 406). Workers in a canning factory are called “boiler makers” and are typically involved in bending steel plates and joining them with rivets. Workers are divided into rivetter, holder-up, plater, angle-ironsmith, and caulking. The work is done in groups, with the skilled angle-ironsmith and caulking. The “bōsin” of the rivetter gives the instructions as the head of the group. Wages are paid on a piece-rate contract basis. Because of the contracted wage, the cannery workers’ income is unstable; thus, they are paid a fixed wage (daily wage amount) during the off-season.
@@
Smithing is the second most common industry, accounting for 28% of the total. Panel A of Table A.6 reveals that most of the smithing industry is relatively small scale. They make boulders, nuts, rivets, lathe stops, jigs, and spanners for can manufacturing and construction. They were subcontractors who manufactured metal parts to supply the larger factories in Tsukishima (Department of Health, Ministry of the Interior 1923a, p. 412). There are three types of jobs: yokoza, sakite, and tatara-fuki (bellows-blower). The yokoza is self-employed skilled worker (oyakata), while the other two are young unskilled laborers. In small operations, wages were based on a percentage based on voluntary business discussions with the proprietor, and working hours were not regular (pp. 411–413).
@@ B.1 Tsukishima Survey: The First Social Survey in Japan
Kawai (1980) summarizes the history of social research methodology. The methodology of social research in Japan was formed in the 1930s. Social research methodologies had been developed in the 1910s and 1920s (Kawai 1980, pp. 47–48). During this period, “numerous social surveys were conducted by various survey entities, including government agencies, research groups, local administrative bodies, private organizations, survey research institutes, and researchers” (p.48). This is because “around WWI, there were many problems in labor, lifestyle, urban, health and sanitation, public safety, and civic due to the heavy and chemical industrialization, the monopolization of capitalism, and urban expansion.” The Report on Field Survey in Tsukishima, Kyobashi Ward, Tokyo (called Tsukishima Survey) [^79] Takano himself, who was responsible for overseeing the survey, refers to this field survey as the “Tsukishima Survey” (Department of Health, Ministry of the Interior 1923a, p. 6). Considering this, I call the Tsukishima Survey throughout this paper. used in this study is the most pioneering of these surveys of urban communities. Sekiya (1970, p. 43) regards the Tsukishima Survey as a pioneering household survey in Japan.
@@
First, I summarize the investigators, research questions, and subjects of the Tsukishima Survey. The person responsible for the Tsukishima Survey was Iwasaburō Takano (hereafter “Takano”), a professor of social statistics and social policy at the Law School of Tokyo Imperial University. Takano became a professor of economics at Tokyo Imperial University in 1919, and became the director of the OISR in 1920. Takano placed great importance on gathering objective statistics about workers through social surveys (Sekiya 1970, p. 33). He had the experience of conducting the first household survey in Japan in 1916.808080This is called the “Survey of Household Income and Expenditure of 20 Artisans and Workers in Tokyo” and was inspired by Charles Booth (Miyoshi 1980, p. 33). Although the original documents of this survey were preserved at Tokyo Imperial University, those were destroyed in the Great Kantō Earthquake (Takano 1933, pp. 727–728). More precisely, before this household survey, Takano conducted a field survey in a poor area of Tokyo when he was a student at Tokyo Imperial University. The findings were reported in April 1894 as an exercise entitled “East London in Tokyo” (Kawai 1981b, p. 3). Takano has been a member of the seventh subcommittee (Rural Sanitation Subcommittee) of the Ministry of Home Affairs’ Health and Sanitation Investigation Committee since 1916.818181The Health and Sanitation Investigation Committee of the Ministry of Home Affairs was established on June 27, 1916, under the second Shigenobu Okuma Cabinet by issuing Imperial Ordinance No. 172. Initially, 34 members were appointed, and the first meeting was held on July 8. The reason for the establishment of this committee was based on the need for a policy response regarding the establishment and institutionalization of various investigation organizations and the penetration of the capitalist economy during the Meiji period. Another reason is the requirement for a public health response to the rising mortality rate, exemplified by the tuberculosis epidemic (Kawai 1981a, pp. 10–13). In 1917, the Seventh Section evolved into the Section on Urban and Rural Sanitary Conditions (Kawai 1981a, pp. 12–13). On October 22, 1918, Takano submitted a proposal to the subcommittee for a field survey of urban sanitary conditions, which was approved. The survey aimed to determine the housing conditions, household budget status, and health status (child health, production, stillbirths, deaths, and diseases) in the areas where skilled-worker families congregated (Sekiya 1970, p. 6). Initially, the survey area was planned to be Yanagibashi Yokogawa-chō in Honjo Ward. However, at a committee meeting on November 23, the survey site was changed to Tsukishima in Kyobashi Ward, which had more skilled worker households than in Yokogawa-chō (Kawai 1980, pp. 55–59). [^82] Takeda (2015, pp. 75–76) provides more details on how the survey site was changed. Takano was looking for a commissioned researcher to conduct the survey and asked Sakuzō Yoshino, a colleague at Tokyo Imperial University, to recommend a suitable person. Yoshino approached his former student Kotora Tanahashi, who had become a member of the fraternity club, a mutual aid organization for workers, to accept the investigator position. The diary written by Kotora Tanahashi (“Tanahashi Kotora Diary”) is in the possession of the OISR, and Takeda used the diary as the basis for her analysis of the circumstances surrounding the change of survey location. Tanahashi eventually recommended Yoshitsuru Yamana, who was his classmate in high school and Tokyo Imperial University, as a researcher and considered Tsukishima to be the best research area because of the ease of selecting survey households using the fraternity’s network (Takeda 2015, pp. 76–77). On October 26, Tanahashi and Yamana, together with the Kyobashi Branch of the fraternity club, conducted an on-site inspection of Tsukishima. Based on this inspection, Tanahashi and Yamana approached Takano about locating the survey station in Tsukishima. Finally, Takano and Yamana conducted an on-site inspection of Tsukishima on October 30 and decided on Tsukishima as the survey site.
@@
Although the statistical method of sampling was not yet common at that time, Takano selected the survey sites based on his visionary thinking. In fact, “There can be no dispute that we have chosen Tokyo as the target of our metropolitan survey. However, it is impossible to survey the entire Tokyo area. We have no choice but to conduct a partial survey. For this purpose, it would be the best idea to select a representative area in Tokyo and survey that area narrowly but deeply.” (Department of Health, Ministry of the Interior 1923a, pp. 2–3). Takano further explained the need to focus on the working-class, which makes up most of the citizens, rather than to see the political and commercial center of Tokyo: “Most of Tokyo’s citizens are wage earners who subsist on labor. As is usual in large cities, a large number of people are engaged in manual labor in Tokyo. However, the most genuine type of worker is the skilled worker. Therefore, the area where many families of skilled workers congregate can be regarded as a representative area of Tokyo. And it seemed to me that a study of the social and sanitary conditions in that area could explain a large part of the social and sanitary conditions in Tokyo.” (Department of Health, Ministry of the Interior 1923a, pp. 3).
@@
It is necessary to delve into the point where Takano states that “the most genuine type of worker is the skilled workers” and the fact that the main target of the Tsukishima Survey was skilled workers. [^83] In other words, the Tsukishima Survey was not designed for poor households. Therefore, unlike the surveys conducted in the UK by C. Booth and B.S. Rowntree, no standard of living was set in the survey (Sekiya 1970, p. 42). The description by Kawai (1981a, pp. 13–18) is helpful in this regard. As summarized in Section 2, Japan’s heavy and chemical industrialization throughout WWI led to a large concentration of city workers. As a result, the ratio of small-scale conventional factories based on “master employees” as is often observed in the textile industry decreased. On the other hand, the number of factories with a direct “factory owner-employee” employment relationship increased, particularly in the machinery and equipment, metal, and chemical industries. These workers in the modern industrial sector accumulated skills, formed households, and settled in the cities. It can be said that a new urban worker lifestyle was being formed. In short, although the city had a class of workers engaged in traditional industries, a new type of workers in the modern industrial sector, and a lower class, Takano believed that the skilled workers in the modern industrial sector would play a central role in terms of the labor force, labor movement, and urban life.
@@
In addition, Sekiya notes the rise of the labor movement. As noted in Section 2, while wages rose throughout WWI, the price surge caused relative wages to decline. In 1916, Japan’s first labor protection legislation, the Factory Law, was enacted. However, the number of working days and hours did not decrease, and the number of labor disputes steadily increased (Sekiya 1970, pp. 24–26). Given this, Takano believed that workers should organize labor unions to help themselves with labor problems that could not be resolved through labor protection legislation. Sekiya (1970, pp. 26–40) speculates that Takano’s intention in conducting the Tsukishima Survey was to obtain basic data for solving labor problems, specifically to identify the scope of matters to be solved by workers on their own initiative (or covered by labor protection legislation). His insights suggest that this historical background should be addressed as a reason for Takano’s focus on skilled workers.
@@
Kawai (1981b, pp. 15; 17) assessed the 40 households covered by the household survey included in the Tsukishima Survey as “the representativeness of the sample has remained ambiguous.” There is some mention but no clear description of how the sample households were characterized among the worker households in Tsukishima. It is reported that “to find suitable persons to fill in the form, we must ask for the assistance of school principals, police officers, physicians, and laborers, and we have requested a meeting of these persons at the survey office to explain the purpose of the survey, and to solicit applicants for the distribution of the household account book” (Department of Health, Ministry of the Interior 1923a, pp. 8–9). As explained, the target households were not randomly selected because the statistical sampling method was not prevalent. However, it is plausible that Takano designed this household budget survey in the sampling framework to capture the representative households in Tsukishima. As described in Section 3, the THBS households can be regareded as the standard skilled factory worker households in Tsukishima.